# UHD-IQA Benchmark Database: Pushing the Boundaries of Blind Photo Quality Assessment

### v1

@@ 5.1 Analysis
Our dataset comprises 6073 labeled examples, all of which were resized to a fixed width of 3840 pixels. The height was proportionally scaled to maintain the original’s aspect ratio. Thus, some images are shorter than the standard 2160 pixels.
We split the dataset into about 70%, 15%, and 15% for the training, validation, and test sets, respectively. We report the exact number of examples for each split in Fig. 5(a). Starting from the test split, we employ a stratified sampling strategy to achieve a distribution of the MOS that is as close as possible to a uniform one [23]. Consequently, as shown in Fig. 5(b), the test and validation splits contain a wider range of labeled MOS than the training one. This creates a more challenging scenario for the models, as they must generalize to unseen image quality scores to perform well on our dataset. Thus, we argue that models achieving high performance on our dataset will likely be suitable for contemporary high-quality and resolution images.
Split ##\## of examples
@@
Figure 5: (a): number of examples constituting each dataset split. (b): distribution of the labeled MOS of each dataset split.
@@ 5.2 Model performance evaluation
We investigate the performance of several state-of-the-art NR-IQA methods on our dataset. To provide a comprehensive analysis, we consider models based on a broad range of approaches, such as standard supervised learning [19, 23], self-supervised learning [13, 2], and vision &\&& language [21, 1]. For a fair comparison, we train each model on the training split of our dataset using the hyperparameters specified by the authors in the respective paper. In particular, for methods based on self-supervised learning and vision &\&& language, we keep the encoder weights frozen and learn a linear regressor and the prompts [21], respectively.
@@
We employ multiple metrics for performance evaluation, namely Kendall’s Rank Correlation Coefficient (KRCC), Pearson’s Linear Correlation Coefficient (PLCC), Spearman’s Rank-order Correlation Coefficient (SRCC), Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE). We do not apply logistic regression to the predictions before computing PLCC, as the model fitting results are hard to reproduce across different works [7]. Additionally, we assess the computational efficiency of the models by measuring the number of Multiply-Accumulations operations (MACs) required for a forward pass with an input image size of $3840\times 2160$ pixel.
@@
We report the results on the validation and test set in Tab. 2 and Tab. 3, respectively. First, we observe that HyperIQA [19] achieves significantly lower performance than the other baselines. We attribute this outcome to the fact that HyperIQA computes the final quality score as the average of 25 different $224\times 224$ random patches extracted from the input image. In contrast, the other methods take the whole image as input and output a single quality score. Even though the strategy adopted by HyperIQA leads to considerably fewer MACs, it does not take full advantage of the amount of information contained in high-resolution images. Second, we notice that CLIP-based models, namely CLIP-IQA+[21] and QualiCLIP[1], obtain the best results according to the correlation-based metrics (KRCC, PLCC, and SRCC) metrics but fall behind when considering the error-based ones (RMSE and MAE). We suppose this is because CLIP-IQA+ and QualiCLIP compute the quality score based on the cosine similarity between the input image and two antonym prompts. Due to the way CLIP is trained [17], this cosine similarity can span the entire $[0,1]$ range. Therefore, the predicted quality scores cover a broader range than the one of the ground-truth MOS (see Fig. 5(b)), leading to higher absolute errors. Finally, we observe that the self-supervised learning baselines (i.e. CONTRIQUE [13] and ARNIQA [2]) achieve the most balanced performance, as they obtain good results for all the metrics while requiring fewer MACs than CLIP-based models.
@@
Table 2: Evaluation of the performance of the baselines on the validation set. $\uparrow$ means that higher values are better, $\downarrow$ means that lower values are better. Best and second-best scores are highlighted in bold and underlined, respectively.
Method KRCC\>$\uparrow$ PLCC\>$\uparrow$ RMSE\>$\downarrow$ MAE\>$\downarrow$ SRCC\>$\uparrow$ MACs (G)\>$\downarrow$
HyperIQA [19] 0.359 0.182 0.087 0.055 0.524 211
Effnet-2C-MLSP [23] 0.445 0.627 0.060 0.050 0.615 345
CONTRIQUE [13] 0.521 0.712 0.049 0.038 0.716 855
ARNIQA [2] 0.523 0.717 0.050 0.039 0.718 855
CLIP-IQA+ [21] 0.546 0.732 0.108 0.087 0.743 895
QualiCLIP [1] 0.557 0.752 0.079 0.064 0.757 901
Table 3: Evaluation of the performance of the baselines on the test set. $\uparrow$ means that higher values are better, $\downarrow$ means that lower values are better. Best and second-best scores are highlighted in bold and underlined, respectively.
Method KRCC\>$\uparrow$ PLCC\>$\uparrow$ RMSE\>$\downarrow$ MAE\>$\downarrow$ SRCC\>$\uparrow$ MACs (G)\>$\downarrow$
HyperIQA [19] 0.389 0.103 0.118 0.070 0.553 211
Effnet-2C-MLSP [23] 0.491 0.641 0.074 0.059 0.675 345
CONTRIQUE [13] 0.532 0.678 0.073 0.052 0.732 855
ARNIQA [2] 0.544 0.694 0.074 0.052 0.739 855
CLIP-IQA+ [21] 0.551 0.709 0.111 0.089 0.747 895
QualiCLIP [1] 0.570 0.725 0.083 0.066 0.770 901

### v2

@@ 5.1 Analysis
We aimed to create a benchmark dataset split that will be used to measure the performance of machine learning models in a representative and challenging manner. Models achieving high performance on our test set should suit contemporary, high-quality, high-resolution images. This requires the test subset to be the most diverse. We divided the dataset into three subsets: approximately 70% for training, 15% for validation, and 15% for testing. The exact number of images in each subset is shown in Fig. 4(a). First, we selected the test set using a stratified sampling strategy to ensure that the Mean Opinion Scores (MOS) distribution more closely resembles a uniform one [31]. Next, we selected the validation set from the remaining images and followed a similar sampling procedure. Finally, the remaining examples constituted the training set. Consequently, the test and validation sets exhibit a broader range of MOS values than the training set, as illustrated in Fig. 4(b). Note that when referencing a subset or dataset split, we refer to the official splits introduced in this benchmark.
Subset ##\## of images
@@
Figure 4: (a): Number of images in the training, validation and test set. (b): Distribution of the MOS in each subset.
5.1.1 Measures of Annotation Reliability
While the images were rated on a scale of $[1,100]$, the following calculations use a normalized Mean Opinion Score (MOS) that is based on rescaled ratings to $[0,1]$. We compared the MOS from the first and second rounds to assess the experiment’s reliability or repeatability. On the test set, this yielded an SRCC of 0.93 and a Root Mean Square Difference (RMSD) of 0.03, indicating a high consistency between the two annotation rounds.
Another measure of reliability or repeatability of the experiment was the agreement between groups of participants. We sampled equal-sized non-overlapping groups of participants and compared their MOS. The group’s size went up to 5 from a total of 10 participants. This group-wise MOS comparison provides a lower bound for the overall reliability of 10 vs. 10 participants, which is our primary interest. The results for the test set, illustrated in Fig. 5, indicate an expected SRCC of at least 0.87 and an RMSD of at most 0.055 for the 10 vs. 10 participant comparison. The SRCC for the validation set is slightly lower at 0.81 due to the stratified sampling strategy of the subsets; however, the validation RMSD is 0.054, almost the same as for the test set.
5.1.2 Implications for Model Performance
Figure 5: (a): Root Mean Squared Difference (RMSD) between the MOS of different sized groups of participants on the test and validation sets. (b): SRCC between MOS of groups. The error bars show $\pm$1 standard deviation of the performance metrics, and the dots represent the averages. 200 samples of pairs of groups of non-overlapping participants were randomly drawn to compute the statistics.
@@ 5.2 Model performance evaluation
We investigate the performance of several state-of-the-art NR-IQA methods on our dataset. To provide a comprehensive analysis, we consider models based on a broad range of approaches, such as standard supervised learning [25, 31], self-supervised learning [18, 2], and vision &\&& language [29, 1]. For a fair comparison, we train each model on the training split of our dataset. In particular, for methods based on self-supervised learning and vision &\&& language, we keep the encoder weights frozen and learn a linear regressor and the prompts [29], respectively. We train Effnet-2C-MLSP using images at $512\times 384$px and $1024\times 768$px resolutions. These are generated by down-scaling $3840\times 2160$ crops from our dataset, which are symmetrically zero-padded for those images with insufficient height, which replicates the training approach from the original paper. For the other models, we use (zero-padded) square patches of size $2160$px, except for HyperIQA [25], which employs $224$px crops.
@@
We employ multiple metrics for performance evaluation, namely Kendall’s Rank Correlation Coefficient (KRCC), Pearson’s Linear Correlation Coefficient (PLCC), Spearman’s Rank-order Correlation Coefficient (SRCC), Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE). We do not apply logistic regression to the predictions before computing the PLCC. Additionally, we assess the computational efficiency of the models by measuring the number of Multiply-Accumulations operations (MACs) [^2] We rely on the ptflops library for measuring the MACs. required for a forward pass with an input image size of $3840\times 2160$ pixel.
We report the results on the validation and test set in Tab. 2 and Tab. 3, respectively. Based on their performance metrics, the overall method ranking is similar between the validation and test subsets. However, due to the sampling strategy, the test set contains a wider MOS range, leading to higher correlations.
@@
Regarding individual methods, first, we observe that HyperIQA [25] achieves significantly lower performance than the other baselines. We attribute this outcome to the fact that HyperIQA computes the final quality score as the average of 25 different $224\times 224$ random patches extracted from the input image. In contrast, the other methods take the whole image as input and output a single quality score. Even though the strategy adopted by HyperIQA leads to considerably fewer MACs, it does not take full advantage of the amount of information contained in high-resolution images. Effnet-2C-MLSP aggregates features from downscaled inputs at different resolutions using a two-stream MLSP extractor backbone. This results in predictions that are more correlated with the ground truth than in the case of HyperIQA. We assume that the high-resolution inputs impose a large domain shift from the ImageNet pretraining. Thus, we expect the model to perform better with additional training data. Furthermore, we notice that CLIP-based models, namely CLIP-IQA+[29] and QualiCLIP[1], obtain the best results according to the correlation-based metrics (KRCC, PLCC, and SRCC) metrics but fall behind when considering the error-based ones (RMSE and MAE). We suppose this is because CLIP-IQA+ and QualiCLIP compute the quality score based on the cosine similarity between the input image and two antonym prompts. Due to the way CLIP is trained [23], this cosine similarity can span the entire $[0,1]$ range. Therefore, the predicted quality scores cover a broader range than the one of the ground-truth MOS (see Fig. 4(b)), leading to higher absolute errors. Finally, we observe that the self-supervised learning baselines (i.e. CONTRIQUE [18] and ARNIQA [2]) achieve the most balanced performance, as they obtain good results for all the metrics while requiring fewer MACs than CLIP-based models.
@@
Table 2: Evaluation of the performance of the baselines on the validation set. $\uparrow$ means that higher values are better, $\downarrow$ means that lower values are better. The best and second-best scores are highlighted in bold and underlined, respectively.
Method KRCC\>$\uparrow$ PLCC\>$\uparrow$ SRCC\>$\uparrow$ RMSE\>$\downarrow$ MAE\>$\downarrow$ MACs (G)\>$\downarrow$
HyperIQA [25] 0.359 0.182 0.524 0.087 0.055 211
Effnet-2C-MLSP [31] 0.445 0.627 0.615 0.060 0.050 345
CONTRIQUE [18] 0.521 0.712 0.716 0.049 0.038 855
ARNIQA [2] 0.523 0.717 0.718 0.050 0.039 855
CLIP-IQA+ [29] 0.546 0.732 0.743 0.108 0.087 895
QualiCLIP [1] 0.557 0.752 0.757 0.079 0.064 901
Table 3: Evaluation of the performance of the baseline methods on the test set. $\uparrow$ means that higher values are better, $\downarrow$ means that lower values are better. Best and second-best scores are highlighted in bold and underlined, respectively.
Method KRCC\>$\uparrow$ PLCC\>$\uparrow$ SRCC\>$\uparrow$ RMSE\>$\downarrow$ MAE\>$\downarrow$ MACs (G)\>$\downarrow$
HyperIQA [25] 0.389 0.103 0.553 0.118 0.070 211
Effnet-2C-MLSP [31] 0.491 0.641 0.675 0.074 0.059 345
CONTRIQUE [18] 0.532 0.678 0.732 0.073 0.052 855
ARNIQA [2] 0.544 0.694 0.739 0.074 0.052 855
CLIP-IQA+ [29] 0.551 0.709 0.747 0.111 0.089 895
QualiCLIP [1] 0.570 0.725 0.770 0.083 0.066 901